1. (4 pts) Estimate the numeric value of 5% accurately to 6 decimal places, using Newtons Method.
Answer #1


x=547 = 275 let, f2) = 87_5. 0 - 1139 = yder let , do=2. -2, = xo - + (80) : [2n+1 = 2n- flam 7 71120) Tran W = 2- f12) 2 => yeti - 4176) 7071 1258498957 - x6 = x+ - f(24) = X8% 1.25849895 tl () ... the numeric value of 517 accurately to o de cinmal